show that ( 1, 1, 1), ( 0, 1, 1) and ( 0, 1, -1) generate or spanned R^3
1
Expert's answer
2020-11-12T19:08:46-0500

"\vec a _1 =(1,1,1), \vec a_2=(0,1,1), \vec a_3=(0,1,-1)"

We show that the vectors form the basis.

Consider a linear combination of vectors

"\alpha \vec a_1+\beta\vec a_2+\gamma\vec a_3=\vec 0"

and show that the numbers "\alpha,\beta,\gamma" are equal 0.

"\alpha \vec a_1+\beta\vec a_2+\gamma\vec a_3=\vec 0\\
\alpha (1,1,1)+\beta(0,1,1)+\gamma(0,1,-1)=(0,0,0)\\
\begin{matrix}

 \alpha=0\\
 \alpha+\beta+\gamma=0\\
\alpha+\beta-\gamma=0
 \\\end{matrix}\\
 \alpha=\beta=\gamma=0"

Then vectors "\vec a _1 =(1,1,1), \vec a_2=(0,1,1), \vec a_3=(0,1,-1)" form the basis in "R^3"